You are lying 120 ft away from a tree that is 50 feet tall. You look up at the top of the tree. Approximately how far is your hear from the top of the tree in a straight line? : You are lying 120 ft away from a tree that is 50 feet tall. You look up at the top of the tree. Approximately how far is your hear from the top of the tree in a straight line?
Answer by nerdybill(1042)

You are lying 120 ft away from a tree that is 50 feet tall. You look up at the top of the tree. Approximately how far is your hear from the top of the tree in a straight line?
.
Draw a diagram of the problem and label all the information provided.
.
From your diagram, a "right triangle" is formed by "your ear" the "top of the tree" and the "base of the tree". Once you see that, you now realize that you can apply Pythagorean's theorem:
.
Let x = distance between your ears and the top of tree
then
x^2 = 50^2 + 120^2
x^2 = 2500 + 14400
x^2 = 16900
x = 130 ft